Update 28th Jan.  By Ian Robins.

A brief update on key loco projects as we progress jobs over the winter season.

D1010 - Transmission repairs ongoing. Currently awaiting a visit from the Voith representative with the tools
necessary to split apart the converter rail.  We are also seeking an estimate to perform the body work repairs
so as a 'cost to complete' can be put together.

D7017 - Currently in 'warm storage' pending the procurement of new bogie leaf springs.

D7018 - Miscellaneous tasks ongoing plus attention to the transmission gear changing issue.  The loco is
expected to visit a couple of other railways later this year.

D9526 - Currently undergoing a 250 hour exam. All cylinder heads removed for attention to a sticking valve
issue.

D9518 -  Awaiting a cost estimate to perform the necessary repairs to the horn guides.

D6566 - Available for use by WSR as required.

D6575 - Winter maintenance.  The CO2 fire bottles have recently been changed.

47 077 - Work on the radiator louvers continues. Being currently too heavy for use on the WSR, the way
ahead for the loco is under consideration.